What is the optimal level of output in the short run when TC= Q2 + 10Q+40 ,Q=SO-ASP


1
Expert's answer
2019-11-28T11:52:46-0500

"TC=Q^2+10Q+40" but"Q= SO-ASP"

"TC= (SO-ASP) ^2 +10(SO-ASP) +40"

"TC=SO^2\u2212ASP ^2+10SO\u221210ASP+40"